I arrived this morning and found that filter cavity is misaligned. After some investigation with Magazzu and Vardaro, we conclude that the problem is due to that filter cavity was not locked for a long time and suspended mirrors drift too much away from the good position. The IR beam was put away but brought back by Vardaro so that we can do some test.
The new cal matrix was tested on 4th Aug (logbook is here). This test result is shown in Figure.1. We can see that all the tilts we introduce for filter cavity mirrors can be sensed with good calibration now. We could estimate the error of this sensing but will be done at a later stage.
The new act matrix is measured on 4th Aug (logbook is here). After some calculation based on Figure.2, I implemented the new matrix here (line 68-74). The implementation of the new matrix requires the restart of AFC process, which requires the update of setpoints in the vpm. The details of updating set points are summarized in the attached note. In attached Figure.3, we see the situation of several channels at the moment when AFC process was stopped and recovered. We can see that all channels are brought back to original points and I also put automation back to exec after this operation.
Note that I tried to test the fast shutter for closing SQZ towards FC, but I fotgot to stop automation. This problem was pointed out by Vardaro and summarized in the attached note as well.
The test of new act matrix will be done in the next days.